OpenStudy (anonymous):

Please HELP! I am supposed to get two x and y values but its not happening?! Solve 4xy + 2y2 = 80 and 6xy + 4y2 = 136 for x and y.

OpenStudy (anonymous):

I would solve it like this: 4xy+2*(y^2)=80 6xy+4*(y^2)=136 It's simultaneous equations first of all it's better to eliminate the 'xy' component otherwise it'll get more complicated So scale up your equations equivalently: 24xy+12*(y^2)=480 24xy+16*(y^2)=544 then eliminate 24xy so: 12*(y^2)-480=16*(y^2)=544 solve for y=+/- 4 Then simply sub back into one of your original equations and solve for your two values of x I hope this helps! :)

OpenStudy (anonymous):

thank you so much!